OpenStudy (anonymous):

determine the range of the function f(x)=(x-3)^2-2

OpenStudy (anonymous):

i would start by plugging 0 in for x, what do you get?

OpenStudy (anonymous):

is it -2,3 @billj5

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

its easy

OpenStudy (anonymous):

the answer is -2,3 right @mathmath333

OpenStudy (mathmath333):

for finding the range u have to find the inverse of function that is \(\large\ \begin{align} \color{black}{f'(x)\hspace{.33em}\\~\\}\end{align}\) the domain of \(\large\ \begin{align} \color{black}{f'(x)\hspace{.33em}\\~\\}\end{align}\) will be the range of \(\large \begin{align} \color{black}{f(x)\hspace{.33em}\\~\\}\end{align}\) here \(\large \begin{align} \color{black}{f'(x)=3\pm\sqrt{x+2}\hspace{.33em}\\~\\}\end{align}\) so the \(\large \begin{align} \color{black}{f'(x)\hspace{.33em}\\~\\}\end{align}\) will be positive only for \(\large \begin{align} \color{black}{ x+2>0\hspace{.33em}\\~\\ x>-2\hspace{.33em}\\~\\ }\end{align}\) hence the range of \(\large f(x)\) is \(\large \begin{align} \color{black}{x|x\in \mathbb{R},x>-2\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

so its -2 and what?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

and all real numbers

OpenStudy (anonymous):

thanks so much